OpenStudy (anonymous):

x/y for x =2/3 and y = 6/7

OpenStudy (alexandervonhumboldt2):

2/3*7/6

OpenStudy (alexandervonhumboldt2):

2/3 / 6/7=2/3*7/6

OpenStudy (alexandervonhumboldt2):

14/12=(7*2)/(6*2)=7/6

OpenStudy (anonymous):